OpenStudy (anonymous):

A package contains 9 candy canes, 6 of which are cracked. If 2 are selected, find the probability of getting no cracked candy canes. 0.3852 0.1938 0.2870 0.0833

OpenStudy (raden):

3C2/9C2 = 3/36 = 1/12 = .... ?